Caspian Sea Drinks is considering the production of a diet drink. The expansion of the plant and the purchase of the equipment necessary to produce the diet drink will cost $25.00 million. The plant and equipment will be depreciated over 10 years to a book value of $1.00 million, and sold for that amount in year 10. Net working capital will increase by $1.19 million at the beginning of the project and will be recovered at the end. The new diet drink will produce revenues of $8.78 million per year and cost $2.04 million per year over the 10-year life of the project. Marketing estimates 10.00% of the buyers of the diet drink will be people who will switch from the regular drink. The marginal tax rate is 29.00%. The WACC is 10.00%. Find the IRR (internal rate of return).
Answer #1

Annual depreciation = (cost - salvage value)/10 years = (25-1)/10 = $2.4 million per year

Income statement:

YEARS
1 2 3 4 5 6 7 8 9 10
Revenues 8.78 8.78 8.78 8.78 8.78 8.78 8.78 8.78 8.78 8.78
less: expenses
Annual costs 2.04 2.04 2.04 2.04 2.04 2.04 2.04 2.04 2.04 2.04
Depreciation 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40
EBIT 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34
less: Tax @ 29% of EBIT 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26
Net income 3.08 3.08 3.08 3.08 3.08 3.08 3.08 3.08 3.08 3.08

Next we compute operating cash flow:

YEARS
1 2 3 4 5 6 7 8 9 10
EBIT 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34 4.34
Add: Depreciation 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40 2.40
less: Taxes 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26 1.26
Operating cash flow 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48

As 10% of the buyers of diet drink are those who switch from the regular drink the opportunity cost here will be = 10% of $8.78 million = $0.878 million

Thus the projected cash flows will be:

YEARS
0 1 2 3 4 5 6 7 8 9 10
Operating cash flow 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48 5.48
Changes in Net working capital -1.19 1.19
Capital spending -25.00 1.00
Opportunity cost -0.878 -0.878 -0.878 -0.878 -0.878 -0.878 -0.878 -0.878 -0.878 -0.878
Total cash flow -26.19 4.60 4.60 4.60 4.60 4.60 4.60 4.60 4.60 4.60 6.79

We will now compute the IRR. Note that IRR is the rate which makes NPV as nil.

Year Total cash flow 1+r PVIF PV = cash flow * PVIF
0 -26.19 1.1247 1.00 -26.19
1 4.60 0.89 4.09
2 4.60 0.79 3.64
3 4.60 0.70 3.24
4 4.60 0.62 2.88
5 4.60 0.56 2.56
6 4.60 0.49 2.27
7 4.60 0.44 2.02
8 4.60 0.39 1.80
9 4.60 0.35 1.60
10 6.79 0.31 2.10
NPV 0.0000

Thus IRR = 1.1247 - 1 = 12.47%